Let A = {(x, y, z) E R^3ㅣx - y + z = 0} where E represents "an element of" and R the set of real numbers
A mapping f is defined in R^(2*2) , the set of all square matrices of order 2 with entries in R by f: R^(2*2) → R^2
A → Av where v = (-1 2).
(a) Determine ker F. Give a basis of it and precise its dimension.
(b)Verify that the matrix B = (2 1
4 2)
belongs to ker F, and determine its coordinate vector with respect to the basis you gave in (a) above
1
Expert's answer
2021-02-11T11:43:53-0500

(a) Let us find determine "\ker F" and give a basis of it and precise its dimension.


"\ker F=\big\{\left(\begin{array}{cc} a & b \\ c & d \end{array}\right)\in\mathbb R^{2*2}\ |\ 
\left(\begin{array}{cc} a & b \\ c & d \end{array}\right)\left(\begin{array}{cc} -1 \\ 2 \end{array}\right)=
\left(\begin{array}{cc} 0 \\ 0 \end{array}\right)\big\}
=\big\{\left(\begin{array}{cc} a & b \\ c & d \end{array}\right)\in\mathbb R^{2*2}\ |\ 
\left(\begin{array}{cc} -a+2b
 \\ -c+2d \end{array}\right)=
\left(\begin{array}{cc} 0 \\ 0 \end{array}\right)\big\}=
\big\{\left(\begin{array}{cc} a & b \\ c & d \end{array}\right)\in\mathbb R^{2*2}\ |\ 
a=2b, c=2d\big\}
=\big\{\left(\begin{array}{cc} 2b & b \\ 2d & d \end{array}\right)\ |\ 
b,d\in\mathbb R\big\}"


It follows that "\ker F=\big\{\left(\begin{array}{cc} 2b & b \\ 2d & d \end{array}\right)\ |\ 
b,d\in\mathbb R\big\}=\big\{b\left(\begin{array}{cc} 2 & 1 \\ 0 & 0 \end{array}\right)+d\left(\begin{array}{cc} 0 & 0 \\ 2 & 1 \end{array}\right)\ |\ 
b,d\in\mathbb R\big\}", and therefore the linear independent matrices "\left(\begin{array}{cc} 2 & 1 \\ 0 & 0 \end{array}\right),\left(\begin{array}{cc} 0 & 0 \\ 2 & 1 \end{array}\right)" form a basis of "\ker F." Thus "\dim(\ker F)=2."



(b) Since for the matrix "B =\left(\begin{array}{cc} 2 & 1 \\ 4 & 2 \end{array}\right)" we have that "\left(\begin{array}{cc} 2 & 1 \\ 4 & 2 \end{array}\right)\left(\begin{array}{cc} -1 \\ 2 \end{array}\right)=\left(\begin{array}{cc} 0 \\ 0 \end{array}\right)", "B" belongs to "\ker F". Taking into account that "B=\left(\begin{array}{cc} 2 & 1 \\ 4 & 2 \end{array}\right)=1\cdot\left(\begin{array}{cc} 2 & 1 \\ 0 & 0 \end{array}\right)+2\cdot \left(\begin{array}{cc} 0 & 0 \\ 2 & 1 \end{array}\right)", we conclude that its coordinate vector with respect to the above basis is "(1,2)."